The width of a rectangular room is 5 feet longer than the width. Write and algebraic expression for the perimeter of a room with width x feet.
Answer by stanbon(57361)

The length of a rectangular room is 5 feet longer than the width. Write and algebraic expression for the perimeter of a room with width x feet.
-------------------------------
width = x ft' length = x+5 feet
----------------
Perimeter = 2(length + width)
P = 2(x+5 + x) = 2(2x+5) = 4x + 10 ft
======================
Cheers,
Stan H.